# Artificial Human Intelligence: The role of Humans in the Development of Next Generation AI

### v1

@@ 5.3.4 Texture/Shape Bias and Fine-Tuning in vision models:
An exploration into shape bias reveals that ViT [149], despite exhibiting a higher shape bias, tends to align more closely with human-like error consistency when fine-tuned. In contrast, ResNet maintains a more human-like error distribution. Augmentation techniques or introduction of shape prior could play a pivotal role in reshaping model representations, impacting their similarity to human classification behavior [151].
@@
While Transformer networks, especially ViT, showcase superior accuracy, they also exhibit a higher shape bias, contrary to most ConvNets which predominantly show texture bias [156, 148, 165, 166]. The fine-tuning of models with augmented data further influences their error consistency, with ViT trending towards a more human-like pattern [166]. The anticipated result could stem from the diminished influence of inductive bias. However, a drawback arises with ViTs [167] as they necessitate a significant volume of data for training to achieve satisfactory quality, contrasting sharply with the considerably lower data requirements in humans.
5.3.5 Humans as Multi-Modal Foundation Models:
There are various reasons to consider humans as complex, natural multi-modal foundation models (FMs), particularly when comparing the biological processes involved in human development with the training technique of any multi-modal agent. Human development dictates a specific development order of multiple sensory modalities such as tactile, sound, vision, and vestibular. At early ages, we learn to combine these inputs to form a coherent understanding of the world, just like how multi-modal foundation models (FMs) integrate data from various sources like images, text, and audio. The trend in multi-modal FM training is that different types of data (text, images, audio) are aligned into a shared representation by treating two modalities as the core pair and align the rest of the modalities through additional trainings/fine-tunnings [168]. However, the order in which the sensory inputs are presented to the model training process does not necessarily align with how the brain integrates multiple sensory inputs into a unified understanding of an object or situation. Though humans can definitely serve as unique inspiration for training multi-modal FMs, as biological processes provide insights into optimizing the selection of modality pairs (anchoring, etc.) and determining the sequence of adding modalities during the training phase.
After the critical period (from birth to around the age of two or three), humans begin developing higher levels of predictive abilities and cognitive functions based on early childhood basic learning and gradually fine-tune their knowledge through experience. This is also true with foundation models that undergo pretraining and fine-tuning gradually upon the improved fidelity and availability of a new modality. In humans, early development builds a broad base of general knowledge, perhaps a common sense of causal events or acquire language through general exposure to words and later experiences help refine skills (predictions, grammar, etc.) and adapt to new environments or tasks, much like how models can be fine-tuned for specific application-required tasks after initial training. In fact, this basic learning phase is the fundamental reason for generalizing knowledge learned from one domain to another. We begin to witness similar emergent behaviors in recent development of FMs [169, 170, 75, 122].
The architectural convergence between biological systems and artificial models is also evident between humans and SoTA multi-modal foundation models. In the human brain, different sensory inputs—such as vision, auditory, and haptic—are processed by distinct neural populations or specialized brain regions (like the occipital lobe specialized for vision and the auditory cortex specialized for sound) [63, 113, 91]. Additionally, even within a single demanding sensory modality, such as vision, separate brain regions handle different aspects like spatial frequency, shape, color, and motion. This separation of sensory processing into distinct pathways mirrors the multi-tower architectures of modern foundation models, where different towers (or specialized subnetworks) handle different types of input data, such as text, images, or audio. For instance, in multi-tower models like CLIP (Contrastive Language–Image Pretraining) [171], one tower processes image data while another handles text, and the two are aligned in a shared latent space. Similarly, transformers are increasingly designed to handle multiple modalities with dedicated structures for each, optimizing the way they learn across diverse data types. Just as the brain integrates separate streams of sensory input to form a unified perception of the environment, multi-modal FMs combine these towers to perform complex cross-modal tasks, such as generating descriptions from images [172] or synthesizing speech from text [173]. This convergence of architecture between biological and computational systems points to a future where AI systems might have the capacity to replicate even more human-like processing efficiencies/capabilities to build general intelligence for general-purpose tasks.
6 Final Thoughts and Discussions
@@
Despite this observation, outstanding performance of the computational frameworks led many research groups to heavily rely on the idea that incorporating human expertise into machine learning systems could be undesirable for introducing bias and, to some extent, dishonesty [176]. In fact, they proposed that a sizable portion of the machine learning community today would be greatly aided if common sense could be inferred from large-scale pleathora of research with the least amount of prior information that could be inherited from a human model. This trend might be the reason for the emergence of self-supervised learning and more broadly self-motivated or self-improving ML [177]. This new trend, in a way I suppose, represents the sheer desire of many AI leaders to search for alternative intelligences that could have better shortcuts compared to evolutionary trajectory of human intelligence development for convergence to an equilibrium: reaching a potentially super-human intelligence.
We could also formulate a hypothesis that bio-inspired AI has a better potential to be more ethical, fair and less judgmental. This line of thinking suggests that AI models derived from biological systems might be better equipped to embody ethical behavior, fairness, and reduced bias compared to other AI forms. When making such a claim, we assume that the architectural and functional properties of a complex system have a direct impact on all the useful emergent behaviors it generates. Drawing inspiration from biological systems, these models may align more closely with ethical principles, exhibiting less judgmental behavior. Several reasons support this stance. The foremost is the comprehensibility and interpretability of such models: enhanced interpretability aids in integrating fairness, ethics, and human values. Biologically plausible or bio-inspired models, grounded in well-explored biological systems, benefit from extensive studies that unravel their functional anatomy. This inherent comprehensibility simplifies the interpretation of their roles in various contexts, thereby facilitating the integration of ethical considerations.
@@ 7 Conclusion
The hypothesis that similar structural properties could lead to analogous functional outcomes underscores the potential of bio-inspired AI. The coupling between architectural design and algorithmic functionality in these models suggests a closer alignment with human cognitive processes, offering a more natural integration of ethical, aesthetic, and moral considerations. As the debate continues over the future direction of AI technology, bio-inspired approaches present a promising avenue for bridging the gap between computational efficiency and the human-specific attribute imperatives that must guide the evolution of the efforts for artificial general intelligence.
8 Acknowledgement
The ideas presented in this document stem from a reassessment of my interpretations of the human role in the vast domain of intelligence, informed heavily by local talks at MIT (such as those at the Simons Center for the Social Brain, the Center for Brains, Minds & Machines, etc.), interactions with the faculty at MIT as well as in neighborhood institutions and researchers at vision conferences (VSS, etc.), technical literature readings. In particular, collaborative discussions initiated with experts at the Sinha Lab. (in particular Prof. Pawan Sinha) as well as MIT meseum (in particular Seth Riskin) have enormously enriched my thoughts and understanding of deep and provocative matters of human vision and computational neuroscience. I extend my deepest gratitude and appreciation to all those involved in these enlightening exchanges.

### v2

@@ IV-D4 Texture/Shape Bias and Fine-Tuning
An exploration into shape bias reveals that ViT [148], despite exhibiting a higher shape bias, tends to align more closely with human-like error consistency when fine-tuned. In contrast, ResNet maintains a more human-like error distribution. Augmentation techniques play a pivotal role in reshaping model representations, impacting their similarity to human classification behavior.
@@
While Transformer networks, especially ViT, showcase superior accuracy, they also exhibit a higher shape bias, contrary to most ConvNets which predominantly show texture bias [155, 147, 163]. The fine-tuning of models with augmented data further influences their error consistency, with ViT trending towards a more human-like pattern. The anticipated result could stem from the diminished influence of inductive bias. However, a drawback arises with ViTs [164] as they necessitate a significant volume of data for training to achieve satisfactory quality, contrasting sharply with the considerably lower data requirements in humans.
V Final Thoughts and Discussions
@@
Despite this observation, outstanding performance of the computational frameworks led many research groups to heavily rely on the idea that incorporating human expertise into machine learning systems could be undesirable for introducing bias and, to some extent, dishonesty [167]. In fact, they proposed that a sizable portion of the machine learning community today would be greatly aided if common sense could be inferred from large-scale pleathora of research with the least amount of prior information that could be inherited from a human model. This trend might be the reason for the emergence of self-supervised learning and more broadly self-motivated or self-improving ML [168]. This new trend, in a way I suppose, represents the sheer desire of many AI leaders to search for alternative intelligences that could have better shortcuts compared to evolutionary trajectory of human intelligence development for convergence to an equilibrium: Reaching a level of potentially super-human intelligence.
We could also formulate a hypothesis that bio-inspired AI has a better potential to be more ethical, fair and less judgmental. This line of thinking suggests that AI models derived from biological systems might be better equipped to embody ethical behavior, fairness, and reduced bias compared to other AI forms as part of their emergent behavior. When making such a claim, we assume that the architectural and functional properties of a complex system have a direct impact on all the useful emergent behaviors it generates. Drawing inspiration from biological systems, these models may align more closely with ethical principles, exhibiting less judgmental behavior. Several reasons support this stance. The foremost is the comprehensibility and interpretability of such models: enhanced interpretability aids in integrating fairness, ethics, and human values. Biologically plausible or bio-inspired models, grounded in well-explored biological systems, benefit from extensive studies that unravel their functional anatomy. This inherent comprehensibility simplifies the interpretation of their roles in various contexts, thereby facilitating the integration of ethical considerations.
@@ VI Conclusion
The hypothesis that similar structural properties could lead to analogous functional outcomes underscores the potential of bio-inspired AI. The coupling between architectural design and algorithmic functionality in these models suggests a closer alignment with human cognitive processes, offering a more natural integration of ethical, aesthetic, and moral considerations. As the debate continues over the future direction of AI technology, bio-inspired approaches present a promising avenue for bridging the gap between computational efficiency and the human-specific attribute imperatives that must guide the evolution of the efforts for AGI.